Calculate the % quinine by molar mass in quinine sulfate. Note that quinine sulfate exists in the hydrated form: (quinine)2·H2SO4·2H2O.
Answer #1

Molecular mass of Quinine sulphate dihydrate ( (quinine)2·H2SO4·2H2O) is 782 g/mol

Molecular mass of H = 1 g/mole

Molecular mass of S = 32 g/mole

Molecular mass of O = 16 g/mole

Molecular mass of H2O = (2x1+16) = 18 g/mole

Molecular mass of H2SO4 = (2x1+32+4x16) = 98 g/mole

therefore, from molecular formula, we can calculate the molecular mass of quinine

Let the Mass of quinine be X

(quinine)2·H2SO4·2H2O

2X + 98 + 2 (18) = 782

2X + 98 + 36 = 782

2X = 782 - 134

X = 648

Percentage of quinine in Quinine sulphate dihydrate ( (quinine)2·H2SO4·2H2O)

Mass.of Quinine Mass.of Quininesulphatedihydrate X 100

648 787 X 100

                                

                                      = 82.86 %